If a buffer solution is 0.240 M in a weak base (Kb = 8.9 × 10-5) and 0.420 M in its conjugate acid, what is the pH?
Answer #1

mixture of weak base and its conjugate acid act as basic buffer

pOH = pKb + log [conjugate acid] / [base]

pKb = - log Kb = - log [8.9 x 10-5] = 4.05

pOH = 4.05 + log [0.420] / [0.240]

pOH = 4.30

pH = 14 - 4.30

pH = 9.70